# Automorphisms of relatively hyperbolic groups and the Farrell–Jones Conjecture

### 23 Nov 2023 v1

@@ 8. Torsion in the infinitely ended case
The aim in this section is to prove as much of Section 6 as possible without the assumption that $G$ is virtually torsion free. To this end we give two propositions. The first allows us – in many cases – to pass from an infinitely-ended group to a finite index subgroup which is a free product of one-ended groups and a finitely generated free group, while the second is a generalisation of the sporadic case (of Section 6.B) to graphs of groups with one, finitely stabilised, edge.
Proposition 8.1.
Suppose $G$ is finitely generated, infinitely-ended and accessible. Further suppose that for each one-ended vertex group $G_{v}$ in a Stallings–Dunwoody decomposition of $G$, there is a normal finite index subgroup $K_{v}$ that trivially intersects all the incident edge groups. Then $G$ has a finite index subgroup $K$ that is a free product of one-ended groups and a finitely generated free group.
Furthermore, for any automorphism $\Phi$, the suspension $G_{\Phi}$ has a finite index subgroup $K_{\Psi}$, where $\Psi$ is the restriction of some power of $\Phi$ to $H$.
Note that [AGHK23, Lemma 5.4] is a similar result, but the statement and proof are considerably simplified by the underlying assumption of that paper that all groups considered are residually finite. Here we prove necessary and sufficient conditions on the vertex groups and incident edge groups for the existence of such a subgroup.
Remarks 8.2.
We record a number of observations about the hypotheses.
  • Accessibility is only used to ensure that the vertex groups are one-ended. More generally, given a splitting over finite edge groups and where the vertex groups satisfy the separability hypothesis, one can find a finite index free product where the non-trivial vertex groups are finite index subgroups of (conjugates of) the original infinite vertex groups.
  • The sufficient conditions given in the statement are also necessary: any finite index $K$ satisfying the conclusion will have intersections $K\cap G_{v}$, finite index in $G_{v}$ avoiding the edge groups. Passing to the normal core recovers the finite index normal subgroup having the desired property.
  • Finite generation is only used in the “furthermore”; a sufficient condition on $\Phi$ would be that it has such a power.
The proof involves Bass–Serre covering theory, see [Bas93] for details. Since it is independent of the rest of the paper we do not provide a self-contained description of the theory here, though references to the appropriate results and definitions will be made as necessary.
Proof.
Let ${\mathcal{G}}$ be the graph of groups corresponding to a Stallings–Dunwoody decomposition of $G$. We will produce a graph of groups ${\mathcal{K}}$ which covers ${\mathcal{G}}$ (in the sense of [Bas93, Definition 2.6]). To do this we need to produce a graph with a map to the underlying graph of ${\mathcal{G}}$, and its vertex (and edge, though these will be trivial) groups, together with monomorphisms to the appropriate vertex group of ${\mathcal{G}}$. Abusing notation, we refer to all these graph and group maps as $f$, trusting that it will be clear from context which is meant. For each edge (and inverse edge) $e$ of ${\mathcal{K}}$ we also need to define an element $\delta_{e}\in G_{f(\iota(e))}$. These allow us to assemble the “star maps” at a vertex $v$ of ${\mathcal{K}}$ (see [Bas93, Proposition 2.4]):
$\begin{array}[]{rcl}\bigsqcup_{e\in f^{-1}(e^{\prime}),\iota(e)=v}K_{\iota}(e)&\to&G_{\iota(e^{\prime})}/G_{e}\\ (e,k)&\mapsto&\delta_{e}kG_{e^{\prime}}\end{array}$
which we are required to check are bijective to ensure we defined a covering. (Note that, in general, there are compatibility conditions to check on the $\delta_{e}$. However, these are vacuous when $K_{e}$ are trivial, as they will be for us – see [Bas93, Definition 2.1].)
By [Bas93, Proposition 2.7], once we have defined ${\mathcal{K}}$ and $f$ so that we have a covering, there is an inclusion (defined using $f$) from $H=\pi_{1}({\mathcal{K}})$ to $G=\pi_{1}({\mathcal{G}})$, and the Bass–Serre trees of are the same, with $K$ acting as a subgroup of $G$. In this case, we can use [And23, Lemma 3.4] to calculate the index $[G:K]$: it is the sum (as $x$ ranges over the preimages of any vertex or edge) of the indices $[G_{x}:K_{x}]$. In particular, to get a finite index subgroup, we need to ensure that all these indices are finite, and that the sums agree.
To facilitate this, let $K_{v}\trianglelefteq G_{v}$ be as given in the statement when $G_{v}$ is infinite, and if $G_{v}$ is finite, let $K_{v}$ be trivial. Set $d_{v}=[G_{v}:K_{v}]$, and set $d=\mathrm{lcm}(\{d_{v}\})$. (Observe that for $v=\iota(e)$, by the tower law and second isomorphism theorem, $|G_{e}|=[G_{e}:K_{v}\cap G_{e}]=[K_{v}G_{e}:K_{v}]$ divides $[G_{v}:K_{v}G_{e}]\cdot[K_{v}G_{e}:K_{v}]=[G_{v}:K_{v}]=d_{v}$, so this accounts for all edge groups as well.)
Now let the vertex set of ${\mathcal{K}}$ consist of $d/d_{v}$ preimages of each $v$, each with vertex group $K_{v}$. Let each $f(K_{v})$ be the inclusion into $G_{v}$. For the edge groups, let $v=\iota(e)$ and note that in order for the local action of $K_{v}$ on its star to respect the index sum formula, we must have $d_{v}/|G_{e}|$ preimages of $e$ adjacent to a preimage of $v$. Note that by the tower law argument above, $d_{v}/|G_{e}|=[G_{v}:K_{v}G_{e}]$. Summing across the $d/d_{v}$ preimages of $v$, we will see the expected $d/|G_{e}|$ preimages of $e$ adjacent to preimages of $v$. Exactly the same argument applies to $\tau(e)$, and so for every edge $e$ there is a bijection between the “heads” at preimages of $\iota(e)$ and “tails” at preimages of $\tau(e)$ where we would like to attach a preimage of $e$. Picking any explicit bijection, give ${\mathcal{K}}$ edges joining the indicated heads and tails. Set all edge groups to be trivial.
Setting $d$ to be the least common multiple ensures that the resulting graph is connected whatever choices are made. (If it has some smaller connected component, then summing indices over the orbits within it will provide a smaller common multiple.)
Given a vertex $v$ of ${\mathcal{K}}$ and the set of incident preimages of some $e^{\prime}$, let the $\delta_{e}$ range over a set of coset representatives $G_{v}/K_{v}G_{e}$. We now investigate the star maps.
First, we see they are injective. Suppose that $\delta_{e_{1}}k_{1}G_{e}=\delta_{e_{2}}k_{2}G_{e}$. This (together with normality of $K_{v}$ in $G_{v}$) implies that $\delta_{e_{1}}$ and $\delta_{e_{2}}$ represent the same $K_{v}G_{e}$ coset, and so that $e_{1}=e_{2}$. Cancelling the $\delta_{e_{i}}$, we now have that $k_{1}G_{e}=k_{2}G_{e}$. This implies that $k_{1}^{-1}k_{2}$ is contained in $G_{e}$, but since the intersection $G_{e}\cap K_{v}$ is trivial, they must be equal.
To see surjectivity, consider some coset $gG_{e^{\prime}}$. Let $\delta_{e}$ be the previously determined representative of the coset $gK_{v}G_{e^{\prime}}$. Notice that $\delta_{e}^{-1}g$ is contained in $K_{v}G_{e}$, and so there is some element $g_{e^{\prime}}$ of $G_{e^{\prime}}$ so that $k=\delta_{e}^{-1}gg_{e^{\prime}}$ is contained in $K_{v}$. By construction, the copy of $k$ lying inside the $K_{v}$ associated to the coset $\delta_{e}K_{v}G_{e^{\prime}}$ is mapped to $\delta_{e}\delta_{e}^{-1}gg_{e^{\prime}}G_{e}$ which is $gG_{e}$.
Since ${\mathcal{K}}$ is a finite graph of groups with trivial edge groups, $K$ is a free product of its non-trivial vertex groups and a finitely generated free group. The non-trivial vertex groups are (conjugates of) finite index subgroups of the original vertex groups, so are themselves one-ended, as required. Finally, ${\mathcal{K}}$ covers ${\mathcal{G}}$ and the sums of indices over the preimages of any edge or vertex is $d$, so $K$ has finite index in $G$.
For the final statement, note that a finitely generated group has only finitely many subgroups of a given finite index, and so some power $\Psi=\Phi^{k}$ preserves $K$. The suspension $K_{\Psi}$ is finite index in $G_{\Phi}$ by one final application of the tower law. ∎
Even if this fails, we may still be able to proceed in some cases.
@@
The proof being largely an elaboration of the arguments in Section 6.B, here we indicate the necessary changes and references.
@@
First, we have to argue that $\Phi$ preserves the action on the Bass–Serre tree $T$ for this splitting. This follows from [Lev05], or for the two vertex case already from [For02], which give that these one edge splittings are rigid: the unique reduced tree in their deformation spaces, together with the hypothesis that $\Phi$ preserves the vertex groups.
@@
Note that by work of Wall [Wal67, Lemma 4.1] virtually cyclic groups act on the line and have a unique maximal finite normal subgroup which is the kernel of this action. Since $G_{e}$ surjects onto $\mathbb{Z}$, in fact this is the unique maximal finite subgroup.
@@
We need to adapt the proofs of Lemma 6.2 and Lemma 6.3 to the new situation. In the first case, we assume that $e$ and $w^{-1}t^{j}e=w^{-1}e$ are edges in the same cylinder. Then $\langle t\rangle$ and $\langle w^{-1}tw\rangle$ are finite index subgroups of the respective stabilisers, and again the commensurability relation implies that there are powers $n,m\in\mathbb{N}$ so that
@@
To adapt the proof of Lemma 6.3, make the adjustments of the first paragraph of that proof and then consider the whole intersection $G_{v}\cap G_{v^{\prime}}$. This group fixes edges in two distinct cylinders, so is contained inside some $G_{e}\cap G_{e^{\prime}}$, where this intersection is between two virtually cyclic subgroups that are not commensurable. In particular, this means the intersection is finite; in fact its cardinality is bounded by the size of the unique maximal finite subgroup (in either – they are conjugate). This means the action of $G\rtimes_{\Phi}\mathbb{Z}$ on $T_{c}$ is acylindrical.
To finish the proof, we recall the bipartite nature of $T_{c}$, and observe that vertex stabilisers are either isomorphic to $A\rtimes_{\Phi|_{A}}\mathbb{Z}$ (or the same for $B$ – the original vertex stabilisers), or a cylinder stabiliser $\operatorname{Per}(\Phi)\rtimes_{\Phi}\mathbb{Z}$. The first kind is in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$ by hypothesis; the second by Lemma 2.6. ∎
With this in hand, one can begin to try and run the induction argument of Section 6 on a Stallings–Dunwoody decomposition of a more general infinitely ended group. However, there seems as yet to be no analogy for the relative hyperbolicity argument used in the non-sporadic case, and so the induction will not be able to proceed if at some stage we encounter a maximal periodic “Stallings–Dunwoody type splitting” that has more than one edge, and at least one edge with non-trivial stabiliser.
A proof of the following conjecture, the analogy of 3.2 for general infinite ended groups, should complete the proof of A with no assumption on torsion.

### 02 Mar 2026 v2

@@ 8. Torsion in the infinitely ended case
The aim in this section is to prove as much of Section˜6 as possible without the assumption that $G$ is virtually torsion free. To this end we prove a generalisation of the sporadic case (of Section˜6.B) to graphs of groups with one finitely stabilised edge.
@@
The proof being largely an elaboration of the arguments in Section˜6.B, here we indicate the necessary changes and references.
@@
First, we have to argue that $\Phi$ preserves the action on the Bass–Serre tree $T$ for this splitting. This follows from [39], or for the two vertex case already from [20], which give that these one edge splittings are rigid: the unique reduced tree in their deformation spaces, together with the hypothesis that $\Phi$ preserves the vertex groups.
@@
Note that by work of Wall [49, Lemma 4.1] virtually cyclic groups act on the line and have a unique maximal finite normal subgroup which is the kernel of this action. Since $G_{e}$ surjects onto $\mathbb{Z}$, in fact this is the unique maximal finite subgroup.
@@
We need to adapt the proofs of Lemma˜6.2 and Lemma˜6.3 to the new situation. In the first case, we assume that $e$ and $w^{-1}t^{j}e=w^{-1}e$ are edges in the same cylinder. Then $\langle t\rangle$ and $\langle w^{-1}tw\rangle$ are finite index subgroups of the respective stabilisers, and again the commensurability relation implies that there are powers $n,m\in\mathbb{N}$ so that
@@
To adapt the proof of Lemma˜6.3, make the adjustments of the first paragraph of that proof and then consider the whole intersection $G_{v}\cap G_{v^{\prime}}$. This group fixes edges in two distinct cylinders, so is contained inside some $G_{e}\cap G_{e^{\prime}}$, where this intersection is between two virtually cyclic subgroups that are not commensurable. In particular, this means the intersection is finite; in fact its cardinality is bounded by the size of the unique maximal finite subgroup (in either – they are conjugate). This means the action of $G\rtimes_{\Phi}\mathbb{Z}$ on $T_{c}$ is acylindrical.
To finish the proof, we recall the bipartite nature of $T_{c}$, and observe that vertex stabilisers are either isomorphic to $A\rtimes_{\Phi|_{A}}\mathbb{Z}$ (or the same for $B$ – the original vertex stabilisers), or a cylinder stabiliser $\operatorname{Per}(\Phi)\rtimes_{\Phi}\mathbb{Z}$. The first kind is in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$ by hypothesis; the second by Lemma˜2.6. ∎
With this in hand, one can begin to try and run the induction argument of Section˜6 on a Stallings–Dunwoody decomposition of a more general infinitely ended group. However, there seems as yet to be no analogy for the relative hyperbolicity argument used in the non-sporadic case, and so the induction will not be able to proceed if at some stage we encounter a maximal periodic “Stallings–Dunwoody type splitting” that has more than one edge, and at least one edge with non-trivial stabiliser.
A proof of the following conjecture, the analogy of ˜3.2 for general infinite ended groups, should complete the proof of ˜A with no assumption on torsion.
@@
Statements and declarations
Availability of data and material: Not applicable
Financial interests: The authors declare they have no financial interests.